For one of my labs we have the products (13.6g) H2SO4, 3g) H3BO3, (7.5g) NH4F, and the product NH4BF4 and we're supposed to make a balanced chemical equation and find the theoretical yield of the product. Whenever I did it I got:

3H2SO4 + H3BO3 + 4NH4F ---------> NH4BF4 + 3HSO4- + 3NH4+

With the limiting reagent being H3BO3 and producting a yield of 4.92g, does this seem correct? My friend in the same course got a different answer.
Answer #1

3H2SO4 + H3BO3 + 4NH4F ---------> NH4BF4 + 3HSO4- + 3NH4+

3 moles of H2SO4 produce 1 mole of NH4BF4

3*98g of H2SO4 produce 104.8g of NH4BF4

13.6g of H2So4 produce = 104.8*13.6/3*98   = 4.84g

so H2SO4 is limiting reactant

1 mole of H3BO3 produce 1 mole of NH4BF4

62g of H3BO3 produce 104.8g of NH4BF4

3g of H3BO3 produce = 104.8*3/62   = 5.07g of NH4BF4

4 moles of NH4F produce 1 moles of NH4BF4

4*37g of NH4F produce 104.8g of NH4BF4

7.5g of NH4F produce = 104.8*7.5/4*37 = 5.31 g of NH4BF4

H2SO4 is limiting reactant